A 5.5-kg rifle fires an 11 g bullet with a velocity of 1000 m/s. Question: what is the recoil velocity of the rifle
1
Expert's answer
2020-02-18T09:36:00-0500

From the conservation of momentum:


"m_1v_1+m_2v_2=0\to v_1=-\frac{m_2v_2}{m_1}=-\frac{0.011\cdot 1000}{5.5}=-2m/s"


Answer. "-2m\/s" in the opposite direction to the direction of motion of the bullet.